# Sections of Lagrangian fibrations on holomorphic symplectic manifolds

### v4

@@ 5.1 The Dolbeault class of a section
Definition 5.1: Let $\pi:\;M{\>\longrightarrow\>}X$ be a Lagrangian fibration, and $\sigma:\;X{\>\longrightarrow\>}M$ a section. By [BDV, Proposition 2.1] (see also 5.2 below), $\sigma^{*}\Omega$ is a $(2,0)+(1,1)$ closed form on $X$. Assume that $H^{2}(X)$ is generated by a closed $(1,1)$-form $\omega$. Then $\sigma^{*}\Omega-\lambda\omega$ is exact for some $\lambda\in{\mathbb{C}}$. Let $\alpha$ be a 1-form which satisfies $d\alpha=\sigma^{*}\Omega-\lambda\omega$. Then ¯$\overline{\partial}\alpha^{0,1}=0$. The Dolbeaut class of $\sigma$ is the class of $\alpha^{0,1}$ in $H^{1}(\Lambda^{0,*}X,\overline{\partial})=H^{1}(X,{\cal O}_{X})$.
Proposition 5.2:  Let $\pi:\;M{\>\longrightarrow\>}{\mathbb{C}}P^{n}$ be a Lagrangian fibration, $C\subset{\mathbb{C}}P^{n}$ a smooth curve, $C_{U}$ its neighbourhood and $\sigma_{U}:\;C_{U}{\>\longrightarrow\>}M$ a section. Assume that the Dolbeault class $[\alpha^{0,1}]\in H^{1}(C_{U},{\cal O}_{C_{U}})$ of $\sigma_{U}$ vanishes. Then $\sigma^{*}_{U}\Omega$ restricted to a smaller neighbourhood of $C$ can be extended to a closed $(2,0)+(1,1)$-form on ${\mathbb{C}}P^{n}$.
@@
Definition 5.3: Let $U\subset{\mathbb{C}}P^{n}$ be an open subset such that $H^{2}(U)$ is generated by the Fubini-Study form. Let $\rho$ be closed $(2,0)+(1,1)$-form on $U$, and $\rho=d\alpha+\lambda\omega$. Then the $(0,1)$-Dolbeault class of $\rho$ is $[\alpha^{0,1}]\in H^{0,1}(M)$.
@@
Theorem 5.5: Let $\pi:\;M{\>\longrightarrow\>}{\mathbb{C}}P^{n}$ be a Lagrangian fibration, $C\subset{\mathbb{C}}P^{n}$ a general line, let $C_{U}$ be a neighbourhood of $C$, and $\sigma_{U}:\;C_{U}{\>\longrightarrow\>}M$ a section. Then the Dolbeault class of $\sigma_{U}$ vanishes.
Proof: 5.3.
@@ 5.2 Currents on complex manifolds
Remark 5.12: In the literature, this is sometimes called “$(n-p,n-q)$-currents”; we follow the conventions from [Dem].
@@
Definition 5.14: Let $f:\;X{\>\longrightarrow\>}Y$ be a proper holomorphic map of complex manifolds, $\dim_{\mathbb{C}}X=\dim_{\mathbb{C}}Y+k$, and $\alpha$ a $(p,q)$-current on $X$. Define the pushforward $f_{*}\alpha$ using $\langle f_{*}\alpha,\tau\rangle:=\langle\alpha,f^{*}\tau\rangle$, where $\tau$ is any test-form. Then $f_{*}\alpha$ is a $(p-k,q-k)$-current. One should think of $f_{*}$ as of the fiberwise integration.
@@
Proof: Let $Z:=\operatorname{im}\sigma$, and let $[Z]$ be its integration current, defined by $\langle[Z],\tau\rangle=\int_{Z}\tau$. Clearly, $\sigma^{*}(\Omega)=\pi_{*}(\Omega\wedge[Z])$. Let $\dim_{\mathbb{C}}M=2n$. To prove that $\pi_{*}(\Omega\wedge[Z])$ is of type $(2,0)+(1,1)$ it suffices to show that $\pi_{*}(\Omega\wedge[Z])\wedge\tau=0$ for any $(n,n-2)$-form $\tau$. This would follow if we prove that $\pi^{*}\tau\wedge\Omega=0$. This assertion is clear, because in coordinates $\Omega=\sum dp_{i}\wedge dq_{i}$, where $dp_{i}$ are pullbacks of the coordinates on $X$, and $\tau$ is proportional to $dp_{1}\wedge dp_{2}\wedge...\wedge dp_{n}$.
@@ 5.3 Vanishing of the cohomology class of the Dolbeault current
Theorem 5.18:  Let $\pi:\;M{\>\longrightarrow\>}{\mathbb{C}}P^{n}$ be a Lagrangian fibration, $C\subset{\mathbb{C}}P^{n}$ a general line, $C_{U}$ a neighbourhood of $C$, and $\sigma_{U}:\;C_{U}{\>\longrightarrow\>}M$ a section constructed in 3.1. Then the Dolbeault class of $\sigma_{U}$ vanishes.
Proof. Step 1: Let $\pi:\;M{\>\longrightarrow\>}X$ be a Lagrangian fibration, with $\Omega-\pi^{*}\omega=d\alpha$ exact and $\omega$ a closed $(1,1)$-form on $X$. We generalize the Dolbeault class to any piecewise smooth multiple section $Z\subset M$, which is proper and $k$-sheeted (if we count sheets with orientations) over $X$. This is done as follows: we split $Z$ onto smooth pieces, restrict $\alpha$ to each, transport to $X$ with the appropriate sign and sum together, and then take the $\overline{\partial}$-class of the (0,1)-part of resulting current. This is equivalent to taking the current $\pi_{*}(\alpha\wedge[Z])$, using 5.2 to show that its (0,1)-part is $\overline{\partial}$-closed, and taking its $\overline{\partial}$-class.
@@
Step 3: As shown in 3.1 (step 1 of the proof), the section $\sigma_{C}:\;C{\>\longrightarrow\>}M$ is obtained as a representative of an integral homology class on $M$ intersected with $\pi^{-1}(C)$, hence its homology class can be extended to $M$. This implies that the Dolbeault class of $\sigma_{U}$ vanishes; indeed, a restriction of any class in $H^{0,1}({\mathbb{C}}P^{n})$ to $C_{U}$ vanishes, because $H^{0,1}({\mathbb{C}}P^{n})=0$.
Combining this theorem with 4.3, we obtain the main result of this paper.
@@
Proof. Step 1: In [SV2], it was shown that any degenerate twistor deformation of a hyperkähler manifold is Kähler. Let $C\subset{\mathbb{C}}P^{n}$ be a general line in the base, and $\sigma:\;C{\>\longrightarrow\>}M$ a smooth section constructed in 3.1. We extend it to an open neighbourhood $C_{U}$ of $C$, obtaining a map $\sigma_{U}:\;C_{U}{\>\longrightarrow\>}M_{U}$, where $M_{U}:=\pi^{-1}(C_{U})$. We apply 1.2 to construct a degenerate twistor deformation $M^{\prime}_{U}$ of $M_{U}$ such that the section $\sigma_{U}:\;C_{U}{\>\longrightarrow\>}M_{U}^{\prime}$ is holomorphic. If this degenerate twistor deformation can be extended from $M_{U}^{\prime}{\>\longrightarrow\>}C_{U}$ to $M^{\prime}{\>\longrightarrow\>}{\mathbb{C}}P^{n}$, then 4.3, applied to the holomorphic map $C_{U}{\>\longrightarrow\>}M^{\prime}$, gives a rational section of $\pi$.
Step 2: Let $\eta_{U}:=\sigma_{U}^{*}(\Omega)$. By 5.2, $\eta_{U}$ is of type $(2,0)+(1,1)$. To obtain $M^{\prime}$, we have to extend $\eta_{U}$ to a closed form of type $(2,0)+(1,1)$ on ${\mathbb{C}}P^{n}$. As seen in Subsection 5.1, there is a cohomological obstruction to such an extension, which belongs to the group $H^{1}(C_{U},{\cal O})$. [^1] It is possible to see that this group is infinite-dimensional.
Step 3: In 5.3 we proved that the obstruction class in $H^{1}(C_{U},{\cal O})$ (i.e., “the Dolbeault class of the section $\sigma_{U}$”) vanishes. Now, 5.1 implies that the degenerate twistor deformation $M_{U}^{\prime}$ can be extended to $M^{\prime}{\>\longrightarrow\>}{\mathbb{C}}P^{n}$.

### v5

@@ 6.1 The Dolbeault class of a section
Definition 6.1: Let $\pi:\;M{\>\longrightarrow\>}X$ be a Lagrangian fibration, and $\sigma:\;X{\>\longrightarrow\>}M$ a section. By [BDV, Proposition 2.1] (see also 6.2 below), $\sigma^{*}\Omega$ is a $(2,0)+(1,1)$ closed form on $X$. Assume that $H^{2}(X)$ is generated by a closed $(1,1)$-form $\omega$. Then $\sigma^{*}\Omega-\lambda\omega$ is exact for some $\lambda\in{\mathbb{C}}$. Let $\alpha$ be a 1-form which satisfies $d\alpha=\sigma^{*}\Omega-\lambda\omega$. Then $\overline{\partial}\alpha^{0,1}=0$. The Dolbeaut class of $\sigma$ is the class of $\alpha^{0,1}$ in $H^{1}(\Lambda^{0,*}X,\overline{\partial})=H^{1}(X,{\cal O}_{X})$.
Proposition 6.2: Let $\pi:\;M{\>\longrightarrow\>}{\mathbb{C}}P^{n}$ be a Lagrangian fibration, $C\subset{\mathbb{C}}P^{n}$ a smooth curve, $C_{U}$ its neighbourhood and $\sigma_{U}:\;C_{U}{\>\longrightarrow\>}M$ a section. Assume that the Dolbeault class $\alpha^{0,1}]\in H^{1}(C_{U},{\cal O}_{C_{U}})$ of $\sigma_{U}$ vanishes. Then $\sigma^{*}_{U}\Omega$ restricted to a smaller neighbourhood of $C$ can be extended to a closed $(2,0)+(1,1)$-form on ${\mathbb{C}}P^{n}$.
@@
Definition 6.3: Let $U\subset{\mathbb{C}}P^{n}$ be an open subset such that $H^{2}(U)$ is generated by the Fubini-Study form. Let $\rho$ be closed $(2,0)+(1,1)$-form on $U$, and $\rho=d\alpha+\lambda\omega$. Then the $(0,1)$-Dolbeault class of $\rho$ is $\alpha^{0,1}]\in H^{0,1}(M)$.
@@
Theorem 6.5: Let $\pi:\;M{\>\longrightarrow\>}{\mathbb{C}}P^{n}$ be a Lagrangian fibration, $C\subset{\mathbb{C}}P^{n}$ a general line, let $C_{U}$ be a neighbourhood of $C$, and $\sigma_{U}:\;C_{U}{\>\longrightarrow\>}M$ a smooth section which satisfies 4.1 (ii). Then the Dolbeault class of $\sigma_{U}$ vanishes.
Proof: 6.3.
@@ 6.2 Currents on complex manifolds
Remark 6.12: In the literature, this is sometimes called “(np,nq)(n-p,n-q)-currents”; we follow the conventions from [Dem].
@@
Definition 6.14: Let $f:\;X{\>\longrightarrow\>}Y$ be a proper holomorphic map of complex manifolds, $\dim_{\mathbb{C}}X=\dim_{\mathbb{C}}Y+k$, and $\alpha$ a $(p,q)$-current on $X$. Define the pushforward $f_{*}\alpha$ using $\langle f_{*}\alpha,\tau\rangle:=\langle\alpha,f^{*}\tau\rangle$, where $\tau$ is any test-form. Then $f_{*}\alpha$ is a (pk,qk)(p-k,q-k)-current. One should think of $f_{*}$ as of the fiberwise integration.
@@
Proof: Let $Z:=\operatorname{im}\sigma$, and let $[Z]$ be its integration current, defined by $\langle[Z],\tau\rangle=\int_{Z}\tau$. Clearly, $\sigma^{*}(\Omega)=\pi_{*}(\Omega\wedge[Z])$. Let $\dim_{\mathbb{C}}M=2n$. To prove that $\pi_{*}(\Omega\wedge[Z])$ is of type $(2,0)+(1,1)$ it suffices to show that $\pi_{*}(\Omega\wedge[Z])\wedge\tau=0$ for any (n,n2)(n,n-2)-form $\tau$. This would follow if we prove that $\pi^{*}\tau\wedge\Omega=0$. This assertion is clear, because in coordinates $\Omega=\sum dp_{i}\wedge dq_{i}$, where $dp_{i}$ are pullbacks of the coordinates on $X$, and $\tau$ is proportional to $dp_{1}\wedge dp_{2}\wedge...\wedge dp_{n}$.
@@ 6.3 Vanishing of the cohomology class of the Dolbeault current
Theorem 6.18: Let $\pi:\;M{\>\longrightarrow\>}{\mathbb{C}}P^{n}$ be a Lagrangian fibration, $C\subset{\mathbb{C}}P^{n}$ a general line, $C_{U}$ a neighbourhood of $C$, and $\sigma_{U}:\;C_{U}{\>\longrightarrow\>}M$ a smooth section which satisfies 4.1 (ii). Then the Dolbeault class of $\sigma_{U}$ vanishes.
Proof. Step 1: Let $\pi:\;M{\>\longrightarrow\>}X$ be a Lagrangian fibration, with $\Omega-\pi^{*}\omega=d\alpha$ exact and $\omega$ a closed $(1,1)$-form on $X$. We generalize the Dolbeault class to any piecewise smooth multiple section $Z\subset M$, which is proper and $k$-sheeted (if we count sheets with orientations) over $X$. This is done as follows: we split $Z$ onto smooth pieces, restrict $\alpha$ to each, transport to $X$ with the appropriate sign and sum together, and then take the $\overline{\partial}$-class of the (0,1)-part of resulting current. This is equivalent to taking the current $\pi_{*}(\alpha\wedge[Z])$, using 6.2 to show that its (0,1)-part is $\overline{\partial}$-closed, and taking its $\overline{\partial}$-class.
@@
Step 3: Since the section $\sigma_{C}$ satisfies 4.1 (ii), its homology class is obtained as a cap product of $\pi^{-1}(C)$ with an integer homology class $\alpha\in H_{2n}(M,{\mathbb{Z}})$. This implies that the Dolbeault class of $\sigma_{U}$ vanishes; indeed, a restriction of any class in $H^{0,1}({\mathbb{C}}P^{n})$ to $C_{U}$ vanishes, because $H^{0,1}({\mathbb{C}}P^{n})=0$.
Combining this theorem with 5.3, we obtain the main result of this paper.
@@
Proof. Step 1: In [SV2], it was shown that any degenerate twistor deformation of a hyperkähler manifold is Kähler. Let $C\subset{\mathbb{C}}P^{n}$ be a general line in the base, and $\sigma:\;C{\>\longrightarrow\>}M$ a smooth section constructed in 4.1. We extend it to an open neighbourhood $C_{U}$ of $C$, obtaining a map $\sigma_{U}:\;C_{U}{\>\longrightarrow\>}M_{U}$, where $M_{U}:=\pi^{-1}(C_{U})$. We apply 1.2 to construct a degenerate twistor deformation $M^{\prime}_{U}$ of $M_{U}$ such that the section $\sigma_{U}:\;C_{U}{\>\longrightarrow\>}M_{U}^{\prime}$ is holomorphic. If this degenerate twistor deformation can be extended from $M_{U}^{\prime}{\>\longrightarrow\>}C_{U}$ to $M^{\prime}{\>\longrightarrow\>}{\mathbb{C}}P^{n}$, then 5.3, applied to the holomorphic map $C_{U}{\>\longrightarrow\>}M^{\prime}$, gives a rational section of $\pi$.
Step 2: Let $\eta_{U}:=\sigma_{U}^{*}(\Omega)$. By 6.2, $\eta_{U}$ is of type $(2,0)+(1,1)$. To obtain $M^{\prime}$, we have to extend $\eta_{U}$ to a closed form of type $(2,0)+(1,1)$ on ${\mathbb{C}}P^{n}$. As seen in Subsection 6.1, there is a cohomological obstruction to such an extension, which belongs to the group $H^{1}(C_{U},{\cal O})$. [^1] It is possible to see that this group is infinite-dimensional.
Step 3: In 6.3 we proved that the obstruction class in $H^{1}(C_{U},{\cal O})$ (i.e., “the Dolbeault class of the section $\sigma_{U}$”) vanishes. Now, 6.1 implies that the degenerate twistor deformation $M_{U}^{\prime}$ can be extended to $M^{\prime}{\>\longrightarrow\>}{\mathbb{C}}P^{n}$.